# Higher order linear differential equations for unitary matrix integrals: applications and generalisations

### v2

@@
Abstract
We consider characterisations of the class of unitary matrix integrals $\langle(\det U)^{q}e^{s^{1/2}{\rm Tr}(U+U^{\dagger})}\rangle_{U(l)}$ in terms of a first order matrix linear differential equation for a vector function of size $l+1$, and in terms of a scalar linear differential equation of degree $l+1$. It will be shown that the latter follows from the former. The matrix linear differential equation provides an efficient way to compute the power series expansion of the matrix integrals, which with $q=0$ and $q=l$ are of relevance to the enumeration of longest increasing subsequences for random permutations, and to the question of the moments of the first and second derivative of the Riemann zeta on the critical line respectively. This procedure is compared against that following from known characterisations involving the $\sigma$-Painlevé III second order nonlinear differential equation. We show too that that the natural $\beta$ generalisation of the unitary group integral permits characterisation by the same classes of linear differential equations.
1School of Mathematics and Statistics, The University of Melbourne, Victoria 3010, Australia.     Email: pjforr@unimelb.edu.au
2Department of Mathematics, University of Sussex, Brighton, BN1 9RH, United Kingdom.
Email: weif0831@gmail.com
3 Department of Mathematics, Technical University of Munich, 80290 München, Germany.
Email: bornemann@tum.de
@@ 1 Introduction
1.1 A unitary matrix integral and the longest increasing subsequence problem
@@
$\Big\langle e^{z{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)},$ (1.1)
the same average with a further factor of $(\det U)^{q}$, and their $\beta$ generalisation. In (1.1) $U(l)$ refers to the group of $l\times l$ unitary matrices sampled with Haar measure. This matrix integral first appeared in the theoretical physics literature on two-dimensional Yang-Mills gauge theory as a partition function [43]. However it is not in this context that interest in differential equations for matrix integrals first arose. Rather, it came about due to the realisation that the coefficients $\{T_{l}(N)\}$ say of $z^{2N}/(N!)^{2}$ in the power series of (1.1) (note that this is even in $z$) are for a given $l$ an integer family, e.g. for $l=4$ and starting at $N=1$ [44, sequence A047889]
$1,2,6,24,119,694,4582,33324,261808,2190688,19318688,178108704,\dots$ (1.2)
Thus, according to [46], $T_{l}(N)$ counts the number or permutations of $\{1,2,\dots,N\}$ such that the length of the longest increasing subsequence is less than or equal to $l$ (this number is $N!$ for $N\leq l$ as illustrated in (1.2)). An alternative specification of $T_{l}(N)$, used in e.g. [10], is as the number of pairs of standard tableaux of content $N$ with length of the first row less than or equal to $l$. For an explanation of these notions, and the equivalence of the two descriptions, see e.g. [34, 47].
In a different guise a formula equivalent to
$\Big\langle e^{z{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}=1+\sum_{N=1}^{\infty}{T_{l}(N)\over(N!)^{2}}z^{2N}$ (1.3)
@@
$\det\Big[I_{j-k}(2z)\Big]_{j,k=1,\dots,l}=1+\sum_{N=1}^{\infty}{T_{l}(N)\over(N!)^{2}}z^{2N},$ (1.4)
@@
where $I_{\nu}(x)={1\over 2\pi}\int_{0}^{2\pi}e^{x\cos\theta}e^{i\nu\theta}\,d\theta$, $\nu\in\mathbb{Z}$, is the modified Bessel function of the first kind (i.e. I-Bessel function). The reason why the Toeplitz determinant in (1.4) is equivalent to the matrix integral in (1.3) follows from the general identities
@@
$\Big\langle\prod_{j=1}^{l}a(\theta_{j})\Big\rangle_{U(l)}={1\over(2\pi)^{l}l!}\int_{0}^{2\pi}d\theta_{1}\cdots\int_{0}^{2\pi}d\theta_{l}\,\prod_{j=1}^{l}a(\theta_{j})\prod_{1\leq j<k\leq l}|e^{i\theta_{k}}-e^{i\theta_{j}}|^{2}\\ =\det\Big[{1\over 2\pi}\int_{0}^{2\pi}a(\theta)e^{i(j-k)\theta}\,d\theta\Big]_{j,k=1,\dots,l}.$ (1.5)
@@
Here the first equality follows from Weyl’s result on the explicit functional form of the eigenvalue probability density function (PDF) for matrices from $U(l)$ with Haar measure [54], [21], while the second may be regarded as a circular analogue of Andréief’s identity for the integral over the product of two determinants [25], with use too of the Vandermonde determinant formula to rewrite the product over the differences squared in this form.
@@
Arguments were given in [35] that the generating function in (1.3) is $D$-finite, meaning that it must satisfy a linear differential equation with polynomial coefficients; for more on this see the first subsection of Appendix B. Later, in [10], these differential equations where determined explicitly for $l\leq 7$. For example, in the case $l=4$ and thus relating to (1.2) one has that (1.3) with $z\mapsto\sqrt{z}$ is a solution of the fifth order equation [10, last collection of displayed equations]
$z^{4}Y^{(5)}+20z^{3}Y^{(4)}-2(10z-59)z^{2}Y^{(3)}-2(91z-110)zY^{\prime\prime}+4(16z^{2}-87z+20)Y^{\prime}+16(8z-5)Y=0$ (1.6)
@@
(there is also a known linear differential equation of degree four for the power series formed according to the rule that the coefficient of $z^{k}$ is given by the $k$-th member of (1.2), and the coefficient of $z^{0}$ is unity [14, §3.3]). In the present paper, using the theory of Selberg correlation integrals [19], [2], [23, §4.6], [31], we will show how to systematically obtain a first order $(l+1)\times(l+1)$ matrix linear differential equation, which has as the first component of its vector solution equal to (1.3). As a preliminary step, the matrix average is expressed in terms of a certain Jack polynomial based hypergeometric function of $l$ variables, specialised to the case that all the variables are equal; see Section 2. For a closely related class of hypergeometric functions, earlier work [28] has identified a matrix differential equation characterisation. It is shown in Section 3 how the Jack polynomial based hypergeometric functions of present interest can be deduced as confluent limits of those appearing in [28], allowing for these too to be similarly characterised. This implies a scalar linear differential equation of degree $l+1$ satisfied by (1.3) (Proposition 4.1 below), which with the use of computer algebra can be made explicit. As an illustration, we do this explicitly for $l=8$, which is the first case beyond the listing in [10]; see Proposition 4.2 below.
@@
Efficient methods for the computation of $\{T_{l}(N)\}$ have come to the fore in the recent works [29, 12] (see too [13]), where a primary aim was to study a scaling regime in the neighbourhood of $l=2\sqrt{N}$ for $N$ large. A precise limit law for the corresponding scaled distribution has been known since [7], and these recent works set about quantifying correction terms. With the numerical tabulations achieved in these works, clear functional forms of corrections were able to exhibited graphically. Such computations were carried out using the known relation for (1.1) and a particular $\sigma$-Painlevé III equation [53, 32], and also its Chazy-I equivalent form. In the second subsection of Appendix B a comparitive discussion of the computational complexity associated with the latter characterisation, and that based on the present matrix differential equation characterisation via the vector recurrence (3.15), is given.
1.2 A unitary matrix integral and the moments of the second derivative of the Riemann zeta function on the critical line
@@
$\Big\langle(\det U)^{l}e^{z{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}=\det\Big[I_{j-k+l}(2z)\Big]_{j,k=0,\dots,l-1}\\ =(-1)^{l(l-1)/2}\det\Big[I_{j+k+1}(2z)\Big]_{j,k=0,\dots,l-1}$ (1.7)
in the recent studies [38], [39]. Here the second equality views the determinant in the Hankel class (i.e. indices a function of $j+k$), which is the structure emphasised in these studies. The theme of [38], [39] is the question of statistical properties, particularly moments, of higher order derivatives of the Riemann zeta function on the critical line. Fundamentally, in keeping with the celebrated work [37], it is hypothesised that these zeta function moments coincide with the scaled limiting moments of the corresponding derivative of the characteristic polynomial of a Haar distributed random unitary matrix; see [36, 18, 20, 55, 9, 8, 3, 6, 26] for earlier work in this direction, and [4, 50, 1, 5, 16] for subsequent developments. In particular, in relation to the second derivative, this leads to the study of [38]
@@
$Z_{N}(s):=e^{-\pi iN/2}e^{i\sum_{n=1}^{N}\theta_{n}/2}s^{-N/2}\prod_{j=1}^{N}(1-se^{-i\theta_{j}}).$ (1.9)
@@
$F_{2}(l)=(-1)^{\frac{l(l-1)}{2}}\sum_{m=0}^{2l}\binom{2l}{m}\left(\frac{d}{dx}\right)^{4l-2m}\left(e^{-\frac{x}{2}}x^{-m-\frac{l^{2}}{2}}f_{m}(x)\right)\Bigg|_{x=0},$ (1.10)
@@
$\det[I_{j+k+1+2\mu_{j+1}}(2\sqrt{x})]_{j,k=0,\ldots,l-1},$ (1.11)
@@
In [39], certain reasoning is given that leads to the conclusion that like (1.1), $\tau_{l}(x)$ satisfies a linear differential equation of degree $l+1$ (see Appendix A for an overview). As mentioned in the opening paragraph, our approach to systematically deriving a linear differential equation that (1.1) satisfies allows for an extra factor $(\det U)^{q}$ to be included for general $q$, and thus includes (1.7). Thus in this regards our work complements [39]. We consider this further in §4.3, and discuss too the advantage of the linear differential equation characterisation over an alternative nonlinear differential equation characterisation put forward in [33].
@@ 2 Jack polynomial based hypergeometric function form
It has been revised in the Introduction that the matrix integral (1.1) has the interpretations as a partition function for a two-dimensional gauge theory, and as the generating function for counting longest increasing subsequences. There is also an occurrence of (1.1) as a limiting gap probability at the hard edge (neighbourhood of the origin) of a positive definite matrix ensemble with unitary symmetry. To specify this, define complex Wishart matrices $W=X^{\dagger}X$ with $X$ an $n\times N$ $(n\geq N)$ standard complex Gaussian matrix (also referred to as a rectangular GinUE matrix [17]). Let $E_{N}((0,s);l)$ with l=nNl=n-N denote the probability of no eigenvalues in $(0,s)$ in this class of complex Wishart ensembles. In [22] it was established that
$E_{\beta=2}^{\rm hard}((0,s);l):=\lim_{N\to\infty}E_{N}((0,s/(4N));l)=e^{-s/4}\Big\langle e^{(s^{1/2}/2){\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}.$ (2.1)
@@
(Here the use of the symbol $\beta$ as a subscript on the LHS is in the context of the underlying Dyson index; see (2.10) below and subsequent text.) The most significant contribution in [22] for present purposes was the identification of the RHS of (2.1) (and generalisations) as a Jack polynomial based generalised hypergeometric function.
@@
To specify this class of multivariable special function, we first recall that, with $\mathbf{x}=(x_{1},\dots,x_{n})$ the Jack polynomials $P_{\kappa}^{(\alpha)}(\mathbf{x})$ are symmetric polynomials of $n$ variables [42, Ch. VI.6], [23, Ch. 12], [41, Ch. 7]. They are dependent on a parameter $\alpha>0$ and labelled by a partition $\kappa:=(\kappa_{1},\dots,\kappa_{n})$; in the latter each part $\kappa_{i}\in\mathbb{Z}_{\geq 0}$ is ordered as $\kappa_{1}\geq\kappa_{2}\geq\cdots\geq\kappa_{n}$. These polynomials are eigenfunctions of the differential operator
$\sum_{j=1}^{n}x_{j}^{2}{\partial^{2}\over\partial x_{j}^{2}}+{2\over\alpha}\sum_{1\leq j<k\leq n}{1\over x_{j}-x_{k}}\Big(x_{j}^{2}{\partial\over\partial x_{j}}-x_{k}^{2}{\partial\over\partial x_{k}}\Big).$ (2.2)
@@
Introducing the dominance partial ordering on partitions of the same length $\mu<\kappa$, the Jack polynomial $P_{\kappa}^{(\alpha)}(\mathbf{x})$ is uniquely determined as the eigenfunction of (2.2) having the triangular structure with respect to the monomial basis $\{m_{\mu}(\mathbf{x})\}$,
$P_{\kappa}^{(\alpha)}(\mathbf{x})=m_{\kappa}(\mathbf{x})+\sum_{\mu<\kappa}c_{\kappa,\mu}^{(\alpha)}m_{\mu}(\mathbf{x}).$ (2.3)
@@
In the definition of Jack polynomial based hypergeometric functions, the Jack polynomials play the role of the monomials, although these must be scaled
$P_{\kappa}^{(\alpha)}(\mathbf{x})\mapsto\alpha^{|\kappa|}P_{\kappa}^{(\alpha)}(\mathbf{x})/h_{\kappa}^{\prime}.$
Here the precise definition of $h_{\kappa}^{\prime}$ (which depends too on $\alpha$) can be found in e.g. [23, Eq. (12.46)]. Playing the role of the Pochhammer symbol is the quantity
$\quad[u]_{\kappa}^{(\alpha)}:=\prod_{j=1}^{n}{\Gamma(u-(j-1)/\alpha+\kappa_{j})\over\Gamma(u-(j-1)/\alpha)};$ (2.4)
@@
see e.g. [23, Eq. (12.46)]. With this notation we then have as the series definition of the generalised hypergeometric function in question (see e.g. [23, §13.1])
@@
${\vphantom{F}}_{p}^{\mathstrut}F_{q}^{(\alpha)}(a_{1},\dots,a_{p};b_{1},\dots,b_{q};\mathbf{x}):=\sum_{\kappa}{\alpha^{|\kappa|}\over h_{\kappa}^{\prime}}\frac{[a_{1}]^{(\alpha)}_{\kappa}\dots[a_{p}]^{(\alpha)}_{\kappa}}{[b_{1}]^{(\alpha)}_{\kappa}\dots[b_{q}]^{(\alpha)}_{\kappa}}P_{\kappa}^{(\alpha)}(\mathbf{x}).$ (2.5)
@@
In the case $n=1$ this is independent of $\alpha$ and reduces to its classical scalar counterpart $_{p}F_{q}$.
@@
We are now in a position to specify the result of [22] which allows the RHS of (2.1) to be identified as a Jack polynomial based generalised hypergeometric function. Thus, for general $\alpha>0$,
@@
${\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}(c+(n-1)/\alpha;(t)^{n})=B_{n}(c,\alpha)\Big({1\over t}\Big)^{(c-1)n/2}$
$\qquad\times\Big({1\over 2\pi}\Big)^{n}\int_{[-\pi,\pi]^{n}}\prod_{j=1}^{n}e^{2t^{1/2}\cos\theta_{j}}e^{i(c-1)\theta_{j}}\prod_{1\leq j<k\leq n}|e^{i\theta_{k}}-e^{i\theta_{j}}|^{2/\alpha}d\theta_{1}\cdots d\theta_{n},$ (2.6)
@@
(see too [23, Eq. (13.27)] and [27, Eq. (4.32)]), where $c\in\mathbb{Z}_{\geq 1}$ (this restriction can be removed by suitably deforming the implied circular contours in (2) to Hankel loops [24, Proof of Prop. 2]) and
$B_{n}(c,\alpha)=\prod_{j=1}^{n}{\Gamma(1+1/\alpha)\Gamma(c+(j-1)/\alpha)\over\Gamma(1+j/\alpha)}.$ (2.7)
@@
The notation $(t)^{n}$ on the LHS denotes that $\mathbf{x}=(t,t,\dots,t)$, and thus all $n$ components are equal to $t$. Making use of (1.5), it follows from (2) that
@@
$\Big\langle e^{s^{1/2}{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}(l;(s)^{l})\Big|_{\alpha=1};$ (2.8)

### v3

# Higher-Order Linear Differential Equations for Unitary Matrix Integrals

@@
[…]
Higher-Order Linear Differential Equations for Unitary Matrix Integrals
[…]
Higher-Order Linear Differential Equations
for Unitary Matrix Integrals:
Applications and Generalisations
(with an Appendix by Folkmar Bornemann)
[…]
Peter J. FORRESTER a and Fei WEI b
[…]
P.J. Forrester and F. Wei
[…]
Received September 04, 2025, in final form January 26, 2026; Published online February 18, 2026
[…]
In this paper, we consider characterisations of the class of unitary matrix integrals $\big\langle(\det U)^{q}{\rm e}^{s^{1/2}\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}$ in terms of a first-order matrix linear differential equation for a vector function of size $l+1$, and in terms of a scalar linear differential equation of degree l+1{l+1}. It will be shown that the latter follows from the former. The matrix linear differential equation provides an efficient way to compute the power series expansion of the matrix integrals, which with $q=0$ and $q=l$ are of relevance to the enumeration of longest increasing subsequences for random permutations, and to the question of the moments of the first and second derivative of the Riemann zeta function on the critical line, respectively. This procedure is compared against that following from known characterisations involving the $\sigma$-Painlevé III second-order nonlinear differential equation. We show too that the natural $\beta$ generalisation of the unitary group integral permits characterisation by the same classes of linear differential equations.
[…]
unitary matrix integral; matrix linear differential equation; random permutation; Riemann zeta function; Painlevé equation
[…]
15B52; 37K10; 05A15
@@ 1 Introduction
1.1 A unitary matrix integral and the longest increasing
subsequence problem
@@ 1.1 A unitary matrix integral and the longest increasing
$\big\langle{\rm e}^{z\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)},$ (1.1)
the same average with a further factor of $(\det U)^{q}$, and their $\beta$ generalisation. In (1.1), $U(l)$ refers to the group of $l\times l$ unitary matrices sampled with Haar measure. This matrix integral first appeared in the theoretical physics literature on two-dimensional Yang–Mills gauge theory as a partition function [43]. However, it is not in this context that interest in differential equations for matrix integrals first arose. Rather, it came about due to the realisation that the coefficients $\{T_{l}(N)\}$ say of $z^{2N}/(N!)^{2}$ in the power series of (1.1) (note that this is even in $z$) are for a given $l$ an integer family, e.g., for $l=4$ and starting at $N=1$ [44, sequence A047889]
$1,2,6,24,119,694,4582,33324,261808,2190688,19318688,178108704,\dots\,.$ (1.2)
Thus, according to [46], $T_{l}(N)$ counts the number or permutations of $\{1,2,\dots,N\}$ such that the length of the longest increasing subsequence is less than or equal to $l$ (this number is $N!$ for $N\leq l$ as illustrated in (1.2)). An alternative specification of $T_{l}(N)$, used in, e.g., [11], is as the number of pairs of standard tableaux of content $N$ with length of the first row less than or equal to $l$. For an explanation of these notions, and the equivalence of the two descriptions, see, e.g., [34, 47].
In a different guise, a formula equivalent to
$\big\langle{\rm e}^{z\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}=1+\sum_{N=1}^{\infty}\frac{T_{l}(N)}{(N!)^{2}}z^{2N}$ (1.3)
@@
$\det[I_{j-k}(2z)]_{j,k=1,\dots,l}=1+\sum_{N=1}^{\infty}\frac{T_{l}(N)}{(N!)^{2}}z^{2N},$ (1.4)
@@
where $I_{\nu}(x)={1\over 2\pi}\int_{0}^{2\pi}{\rm e}^{x\cos\theta}{\rm e}^{{\rm i}\nu\theta}{\rm d}\theta$, $\nu\in\mathbb{Z}$, is the modified Bessel function of the first kind (i.e., the I-Bessel function). The reason why the Toeplitz determinant in (1.4) is equivalent to the matrix integral in (1.3) follows from the general identities
@@
$\left\langle\prod_{j=1}^{l}a(\theta_{j})\right\rangle_{U(l)}$ $={1\over(2\pi)^{l}l!}\int_{0}^{2\pi}{\rm d}\theta_{1}\cdots\int_{0}^{2\pi}{\rm d}\theta_{l}\prod_{j=1}^{l}a(\theta_{j})\prod_{1\leq j<k\leq l}\big|{\rm e}^{{\rm i}\theta_{k}}-{\rm e}^{{\rm i}\theta_{j}}\big|^{2}$
$=\det\left[{1\over 2\pi}\int_{0}^{2\pi}a(\theta){\rm e}^{{\rm i}(j-k)\theta}{\rm d}\theta\right]_{j,k=1,\dots,l}.$ (1.5)
@@
Here the first equality follows from Weyl’s result on the explicit functional form of the eigenvalue probability density function (PDF) for matrices from $U(l)$ with Haar measure [21, 54], while the second may be regarded as a circular analogue of Andréief’s identity for the integral over the product of two determinants [26], with use too of the Vandermonde determinant formula to rewrite the product over the differences squared in this form.
@@
Arguments were given in [35] that the generating function in (1.3) is $D$-finite, meaning that it must satisfy a linear differential equation with polynomial coefficients; for more on this see Appendix B.1. Later, in [11], these differential equations where determined explicitly for $l\leq 7$. For example, in the case $l=4$ and thus relating to (1.2), one has that (1.3) with $z\mapsto\sqrt{z}$ is a solution of the fifth-order equation [11, last collection of displayed equations]
$z^{4}Y^{(5)}+20z^{3}Y^{(4)}-2(10z-59)z^{2}Y^{(3)}-2(91z-110)zY^{\prime\prime}$
$\qquad+4(16z^{2}-87z+20)Y^{\prime}+16(8z-5)Y=0$ (1.6)
@@
(there is also a known linear differential equation of degree four for the power series formed according to the rule that the coefficient of $z^{k}$ is given by the $k$-th member of (1.2), and the coefficient of $z^{0}$ is unity [14, Section 3.3]). In the present paper, using the theory of Selberg correlation integrals (see [2, 19, 31] and [24, Section 4.6]), we will show how to systematically obtain a first-order $(l+1)\times(l+1)$ matrix linear differential equation, which has as the first component of its vector solution equal to (1.3). As a preliminary step, the matrix average is expressed in terms of a certain Jack polynomial based hypergeometric function of $l$ variables, specialised to the case that all the variables are equal; see Section 2. For a closely related class of hypergeometric functions, earlier work [28] has identified a matrix differential equation characterisation. It is shown in Section 3 how the Jack polynomial based hypergeometric functions of present interest can be deduced as confluent limits of those appearing in [28], allowing for these too to be similarly characterised. This implies a scalar linear differential equation of degree $l+1$ satisfied by (1.3) (Proposition 4.1 below), which with the use of computer algebra can be made explicit. As an illustration, we do this explicitly for $l=8$, which is the first case beyond the listing in [11]; see Proposition 4.2 below.
@@
Efficient methods for the computation of $\{T_{l}(N)\}$ have come to the fore in the recent works [13, 29] (see too [12]), where a primary aim was to study a scaling regime in the neighbourhood of $l=2\sqrt{N}$ for $N$ large. A precise limit law for the corresponding scaled distribution has been known since [7], and these recent works set about quantifying correction terms. With the numerical tabulations achieved in these works, clear functional forms of corrections were able to exhibited graphically. Such computations were carried out using the known relation for (1.1) and a particular $\sigma$-Painlevé III equation [32, 53], and also its Chazy-I equivalent form. In Appendix B.2, a comparative discussion of the computational complexity associated with the latter characterisation, and that based on the present matrix differential equation characterisation via the vector recurrence (3.12), is given.
1.2 A unitary matrix integral and the moments of the second derivative
of the Riemann zeta function on the critical line
@@ 1.2 A unitary matrix integral and the moments of the second derivative
$\big\langle(\det U)^{l}{\rm e}^{z\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}$ $=\det[I_{j-k+l}(2z)]_{j,k=0,\dots,l-1}$
$=(-1)^{l(l-1)/2}\det[I_{j+k+1}(2z)]_{j,k=0,\dots,l-1}$ (1.7)
in the recent studies [38, 39]. Here the second equality views the determinant in the Hankel class (i.e., indices a function of $j+k$), which is the structure emphasised in these studies. The theme of [38, 39] is the question of statistical properties, particularly moments, of higher-order derivatives of the Riemann zeta function on the critical line. Fundamentally, in keeping with the celebrated work [37], it is hypothesised that these zeta function moments coincide with the scaled limiting moments of the corresponding derivative of the characteristic polynomial of a Haar distributed random unitary matrix; see [3, 6, 8, 9, 18, 20, 27, 36, 55] for earlier work in this direction, and [1, 4, 5, 16, 50] for subsequent developments. In particular, in relation to the second derivative, this leads to the study of [38]
@@
$Z_{N}(s):={\rm e}^{-\pi{\rm i}N/2}{\rm e}^{{\rm i}\sum_{n=1}^{N}\theta_{n}/2}s^{-N/2}\prod_{j=1}^{N}\bigl(1-s{\rm e}^{-{\rm i}\theta_{j}}\bigr).$
@@
$F_{2}(l)=(-1)^{\frac{l(l-1)}{2}}\sum_{m=0}^{2l}\binom{2l}{m}\left(\frac{\rm d}{{\rm d}x}\right)^{4l-2m}\bigl({\rm e}^{-\frac{x}{2}}x^{-m-\frac{l^{2}}{2}}f_{m}(x)\bigr)\big|_{x=0},$
@@
$\det\big[I_{j+k+1+2\mu_{j+1}}\bigl(2\sqrt{x}\bigr)\big]_{j,k=0,\ldots,l-1},$
@@
In [39], certain reasoning is given that leads to the conclusion that like (1.1), $\tau_{l}(x)$ satisfies a linear differential equation of degree $l+1$ (see Appendix A for an overview). As mentioned in the opening paragraph, our approach to systematically deriving a linear differential equation that (1.1) satisfies allows for an extra factor $(\det U)^{q}$ to be included for general $q$, and thus includes (1.7). Thus in this regards our work complements [39]. We consider this further in Section 4.3, and discuss too the advantage of the linear differential equation characterisation over an alternative nonlinear differential equation characterisation put forward in [33].
@@ 2 Jack polynomial based hypergeometric function form
It has been revised in the introduction that the matrix integral (1.1) has the interpretations as a partition function for a two-dimensional gauge theory, and as the generating function for counting longest increasing subsequences. There is also an occurrence of (1.1) as a limiting gap probability at the hard edge (neighbourhood of the origin) of a positive definite matrix ensemble with unitary symmetry. To specify this, define complex Wishart matrices $W=X^{\dagger}X$ with $X$ an $n\times N$ $(n\geq N)$ standard complex Gaussian matrix (also referred to as a rectangular GinUE matrix [17]). Let $E_{N}((0,s);l)$ with l=nNl=n-N denote the probability of no eigenvalues in $(0,s)$ in this class of complex Wishart ensembles. In [23], it was established that
$E_{\beta=2}^{\rm hard}((0,s);l):=\lim_{N\to\infty}E_{N}((0,s/(4N));l)={\rm e}^{-s/4}\big\langle{\rm e}^{(s^{1/2}/2)\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}.$ (2.1)
@@
(Here the use of the symbol $\beta$ as a subscript on the left-hand side is in the context of the underlying Dyson index; see (2.6) below and subsequent text.) The most significant contribution in [23] for present purposes was the identification of the right-hand side of (2.1) (and generalisations) as a Jack polynomial based generalised hypergeometric function.
@@
To specify this class of multivariable special function, we first recall that, with $\mathbf{x}=(x_{1},\dots,\allowbreak x_{n})$ the Jack polynomials $P_{\kappa}^{(\alpha)}(\mathbf{x})$ are symmetric polynomials of $n$ variables [42, Chapter VI.6], [24, Chapter 12], [41, Chapter 7]. They are dependent on a parameter $\alpha>0$ and labelled by a partition $\kappa:=(\kappa_{1},\dots,\kappa_{n})$; in the latter each part $\kappa_{i}\in\mathbb{Z}_{\geq 0}$ is ordered as $\kappa_{1}\geq\kappa_{2}\geq\cdots\geq\kappa_{n}$. These polynomials are eigenfunctions of the differential operator
$\sum_{j=1}^{n}x_{j}^{2}{\partial^{2}\over\partial x_{j}^{2}}+{2\over\alpha}\sum_{1\leq j<k\leq n}{1\over x_{j}-x_{k}}\left(x_{j}^{2}{\partial\over\partial x_{j}}-x_{k}^{2}{\partial\over\partial x_{k}}\right).$ (2.2)
@@
Introducing the dominance partial ordering on partitions of the same length $\mu<\kappa$, the Jack polynomial $P_{\kappa}^{(\alpha)}(\mathbf{x})$ is uniquely determined as the eigenfunction of (2.2) having the triangular structure with respect to the monomial basis $\{m_{\mu}(\mathbf{x})\}$, $P_{\kappa}^{(\alpha)}(\mathbf{x})=m_{\kappa}(\mathbf{x})+\sum_{\mu<\kappa}c_{\kappa,\mu}^{(\alpha)}m_{\mu}(\mathbf{x})$.
@@
In the definition of Jack polynomial based hypergeometric functions, the Jack polynomials play the role of the monomials, although these must be scaled $P_{\kappa}^{(\alpha)}(\mathbf{x})\mapsto\alpha^{|\kappa|}P_{\kappa}^{(\alpha)}(\mathbf{x})/h_{\kappa}^{\prime}$. Here the precise definition of $h_{\kappa}^{\prime}$ (which depends too on $\alpha$) can be found in, e.g., [24, equation (12.46)]. Playing the role of the Pochhammer symbol is the quantity
$\qquad[u]_{\kappa}^{(\alpha)}:=\prod_{j=1}^{n}{\Gamma(u-(j-1)/\alpha+\kappa_{j})\over\Gamma(u-(j-1)/\alpha)};$
@@
see, e.g., [24, equation (12.46)]. With this notation we then have as the series definition of the generalised hypergeometric function in question (see, e.g., [24, Section 13.1])
@@
${\vphantom{F}}_{p}^{\mathstrut}F_{q}^{(\alpha)}(a_{1},\dots,a_{p};b_{1},\dots,b_{q};\mathbf{x}):=\sum_{\kappa}{\alpha^{|\kappa|}\over h_{\kappa}^{\prime}}\frac{[a_{1}]^{(\alpha)}_{\kappa}\dots[a_{p}]^{(\alpha)}_{\kappa}}{[b_{1}]^{(\alpha)}_{\kappa}\dots[b_{q}]^{(\alpha)}_{\kappa}}P_{\kappa}^{(\alpha)}(\mathbf{x}).$
@@
In the case $n=1$, this is independent of $\alpha$ and reduces to its classical scalar counterpart $_{p}F_{q}$.
@@
We are now in a position to specify the result of [23] which allows the right-hand side of (2.1) to be identified as a Jack polynomial based generalised hypergeometric function. Thus, for general $\alpha>0$},
@@
${\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}(c+(n-1)/\alpha;(t)^{n})={}$ $B_{n}(c,\alpha)\left({1\over t}\right)^{(c-1)n/2}\left({1\over 2\pi}\right)^{n}\int_{[-\pi,\pi]^{n}}\prod_{j=1}^{n}{\rm e}^{2t^{1/2}\cos\theta_{j}}{\rm e}^{{\rm i}(c-1)\theta_{j}}$
$\times\prod_{1\leq j<k\leq n}\big|{\rm e}^{{\rm i}\theta_{k}}-{\rm e}^{{\rm i}\theta_{j}}\big|^{2/\alpha}{\rm d}\theta_{1}\cdots{\rm d}\theta_{n},$ (2.3)
@@
(see too [24, equation (13.27)] and [22, equation (4.32)]), where $c\in\mathbb{Z}_{\geq 1}$ (this restriction can be removed by suitably deforming the implied circular contours in (2.3) to Hankel loops [25, proof of Proposition 2]) and
$B_{n}(c,\alpha)=\prod_{j=1}^{n}{\Gamma(1+1/\alpha)\Gamma(c+(j-1)/\alpha)\over\Gamma(1+j/\alpha)}.$
@@
The notation $(t)^{n}$ on the left-hand side denotes that $\mathbf{x}=(t,t,\dots,t)$, and thus all $n$ components are equal to $t$. Making use of (1.5), it follows from (2.3) that
@@
$\big\langle{\rm e}^{s^{1/2}\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}\bigl(l;(s)^{l}\bigr)\big|_{\alpha=1};$ (2.4)